# Topology of the simplest gene switch

### v1

@@ Abstract
turns out to be $+1$ for all three regimes. When the network ends in a sink, i.e. extinction in the regime of death with strictly no regeneration, however the winding number becomes 0. In this paper, we will focus on ergodic networks that do regenerate, albeit at a small rate.
In the non-adiabatic phase, local winding number predicts bistability and location of steady state peaks.—While the global spectral winding can distinguish between different dynamical regimes and identifies the phase transitions, it would be useful to predict specific quantities of interest within each phase, such as the location of the steady-state peak and the relative heights of different peaks. Here, we introduce a local winding number for this purpose. The presence of local winding numbers in a single one-dimensional chain can predict the accumulation of probability density at the system edge, also known as the non-Hermitian skin effect [41, 42, 29, 43]. However, these winding numbers are usually calculated under periodic boundary conditions [29, 43]. Here, the absence of translation symmetry in transcription factor number in many networks including ours, renders this approach inapplicable.
Figure 4: In the adiabatic phase, combined local winding determines domain wall and peak location. (a) Local winding number is tracked along two averaged chains (black). Cartoons illustrate two networks with periodic boundary conditions, where the unbound (resp. bound) chain dominates the net direction of winding. (b) Steady state peak in the total probability (black) is determined by the domain wall in the local winding number. (c) Observed peak location $n_{p}$ (black dots) coincides with the predicted location of the domain wall $n^{*}_{\mathrm{tot}}$ (lines of shades of gray) for a range of parameter values. (d) Relative probability between unbound and bound chains $\Delta p=\sum_{n}p_{n1}-p_{n0}$ for a range of parameters $X^{\mathrm{ad}}$ and $X^{\mathrm{eq}}$. Probabilities in two chains are equal when $X^{\mathrm{eq}}=X^{\mathrm{ad}}$. Insets: when $X^{\mathrm{eq}}>X^{\mathrm{ad}}$, unbound probability is larger than the bound such that upward and downward fluxes are balanced. Similarly, unbound probability is smaller than the bound when $X^{\mathrm{eq}}<X^{\mathrm{ad}}$. We use $X^{\mathrm{eq}}=X^{\mathrm{ad}}=20$ and $\omega=10$ for plots.
To address the lack of translation symmetry where network parameters are changing as we move along protein number $n$, we develop a local winding number that is evaluated at each index $n$ of the network. In the non-adiabatic phase $\omega\ll 1$, slow transitions between the bound and unbound chains allow the gene switch to be considered as two separate chains. We can thus define a local winding for the upper and lower (bound and unbound respectively) chains separately. For each chain, we choose a unit cell that consists of the $n$-th state of the chain together with transitions connecting it to the right neighbor. Repeating this unit cell many times leads to new periodic boundaries such that the transition matrix of the bound chain becomes $\mathcal{W}^{\mathrm{n0}}(\chi)=(n+1)k(e^{-i\chi}-1)$, and that of the unbound chain becomes $\mathcal{W}^{\mathrm{n1}}(\chi)=g_{1}(e^{i\chi}-1)+(n+1)k(e^{-i\chi}-1)$.
We compute the local winding in both bound and unbound chains by plugging these transition matrices in Eq. (1), to show in Fig. 3(a) how they change with the protein number $n$ along the network. In the bound chain, the local winding is always $\nu_{\mathrm{n0}}=-1$ along the entire chain, suggesting a polarization to the left that leads to the steady state localized on the left edge at $n^{*}_{0}=0$ (blue in Fig. 3(b)). In the unbound chain, the local winding number switches from $\nu_{\mathrm{n1}}=+1$ to $\nu_{\mathrm{n1}}=-1$ at the protein number $n^{*}_{1}=g_{1}/k-1=2X^{\mathrm{ad}}-1$, which is where the forward and backward transitions become equal to each other. This signals a domain wall at $n^{*}_{1}$, where we expect an edge state. Indeed, we find the steady state peak at this domain wall; see the orange line in Fig. 3(b) for $X^{\mathrm{eq}}=X^{\mathrm{ad}}=20$. We verify this analytical prediction by comparing it to the numerical solution of the steady state peak across a range of system parameters $X^{\mathrm{ad}}$ and $X^{\mathrm{eq}}$, to find that they match well (see Fig. 3(c)).
We can further predict which of the two peaks, bound or unbound, will be larger. Here, we use a principle of flux balance similar to Kirchhoff’s laws for electrical circuits. This principle states that the total upward flux in the gene network must be equal to the total downward flux such that $\sum_{n}J_{(n+1)1,n0}=0$. Here, the probability flux along a transition between $i$-th and $j$-th state is defined as $J_{ji}=\mathcal{W}_{ji}p_{i}-\mathcal{W}_{ij}p_{j}$. Since at the steady state there are two narrow peaks in the non-adiabatic phase, we can use flux balance to determine the relative sizes of the peaks. The bound peak at $n=0$ with the total probability $p_{0}$ causes an upwards flux $J_{\uparrow}\approx fp_{0}$. At the same time, the unbound peak at $n\approx 2X^{\mathrm{ad}}$ with the total probability $p_{1}$ causes a downwards flux $J_{\downarrow}\approx 2X^{\mathrm{ad}}hp_{1}$. Since $J_{\uparrow}=J_{\downarrow}$, the two peaks are equal when $2X^{\mathrm{ad}}=f/h=X^{\mathrm{eq}}$ (dashed line in Fig. 3(d)). We predict that the unbound peak is larger than the bound peak when $X^{\mathrm{eq}}>2X^{\mathrm{ad}}$ and vice versa, which we also confirm numerically in Fig. 3(d).
In the adiabatic phase, local winding number predicts mono-stability and location of steady-state.—In the adiabatic phase $\omega\gg 1$, fast transitions between the bound and unbound chains cause the system to average over these states, similar to the Shea-Ackers model [44]. Due to this averaging, we define a combined local winding number for coupled bound and unbound states. This local winding would then predict the steady state of the full gene network.
To define this invariant, we choose a unit cell that consists of the $n$-th state in the bound chain and the $(n+1)$-th state in the unbound chain, together with transitions between these states as well as transitions to and from their right neighbors (see SM for details). Imposing periodic boundary conditions on this unit cell, we obtain the combined local periodic network and compute the corresponding local winding number using Eq. (1). Plotting the combined local winding as a function of the protein number $n$ in Fig. 4(a), we observe that it changes sign over a domain wall, when the winding of the bound chain starts dominating over the unbound chain, as illustrated in cartoons above the plot.
The domain wall in the combined winding number determines the location of the steady state peak, as we illustrate in Fig. 4(b). To analytically derive the domain wall location, we use the fact that the sign of the winding number equals to the sign of the probability flux along the periodic network [29]. Therefore, the local winding number changes sign where probability flux vanishes. We use this to derive that the domain wall (details in the SM) is located at $n^{*}_{\mathrm{tot}}=\left(-X^{\mathrm{eq}}+\sqrt{(X^{\mathrm{eq}})^{2}+4X^{\mathrm{eq}}(2X^{\mathrm{ad}}-1)}\right)/2$. To confirm this prediction, we verify in Fig. 4(c) that the domain wall determines the peak location for a range of parameters $X^{\mathrm{ad}}$ and $X^{\mathrm{eq}}$. We note that when defining the global winding number, we used the same value of protein number $n^{*}=n^{*}_{\mathrm{tot}}$ to introduce the complex phase factor.
Similar to what was done for the non-adiabatic phase, we can use flux balance to predict the probability distribution between bound and unbound chains. With both bound and unbound peaks located at $n=n^{*}_{\mathrm{tot}}$, the up and down fluxes are $J_{\uparrow}\approx fp_{0}$ and $J_{\downarrow}\approx n^{*}_{\mathrm{tot}}hp_{1}$. Probabilities are equally distributed when $n^{*}_{\mathrm{tot}}=f/h=X^{\mathrm{eq}}$, which is true when $X^{\mathrm{ad}}\approx X^{\mathrm{eq}}$ (dashed line in Fig. 3(d)). We numerically confirm this prediction in Fig. 3(d).
Conclusion.—We have developed topological methods for stochastic systems that break translation symmetry and demonstrate them on the simplest gene network. We see there are three different phases even for this simple network, and several crucial properties of its steady state, such as peak location and relative heights, can be understood topologically. These results pave the way for future explorations using the lens of topology. For instance, relaxation times are predicted to scale differently due to topology [29], which would have biological consequences. Another interesting regime is when the global winding number goes to zero, i.e. the non-ergodic limit [43] which applies where the transcription factor can go extinct without the possibility of regeneration.
[…]
@@
Acknowledgments.—We thank Dexin Li and Erwin Frey for helpful insights into winding numbers and flux balance respectively. This work was supported by the NSF Center for Theoretical Biological Physics (PHY-2019745) and the NSF CAREER Award (DMR-2238667).
Supplemental Material to: Topology of the simplest gene switch
Aleksandra Nelson,1 Peter Wolynes,1, 2, 3, 4 and Evelyn Tang1,3
1Center for Theoretical Biological Physics, Rice University, Houston TX, USA
2
Department of Chemistry, Rice University, Houston TX, 77005, USA
3
Department of Physics and Astronomy, Rice University, Houston TX, 77005, USA
4
Department of Biosciences, Rice University, Houston TX, 77005, USA
(Dated: March 3, 2025)
A Exceptional point phase transition
In the main text, we discussed that three regimes of the gene switch are separated by exceptional point phase transitions [39, 38]. Here, we provide a more detailed analysis of these phase transitions.
Exceptional points are characterized by a simultaneous coalescence of eigenvalues and eigenvectors of a non-Hermitian matrix. We demonstrate that the two closest to the steady state eigenvalues of the transition matrix $\mathcal{W}$ go through an exceptional point. For this we plot in Fig. S1 how their real and imaginary parts change with adiabaticity $\omega=f/k$. We also plot a normalized scalar product of the corresponding eigenvectors, which turn to 1 when the two vectors become linearly dependent. This happens at the same values of $\omega$ at which the eigenvalues coalesce, proving that the system goes through exceptional points phase transitions.
Plotting in Fig. S1 the coherence of gene network oscillations for the same parameter range, we confirm that coherence changes from zero to non-zero values at exceptional point phase transitions.
@@ A Exceptional point phase transition
Figure S1: Exceptional point phase transition in the gene switch. We plot two closest to the steady state eigenvalues, and a scalar product of the corresponding eigenvectors as a function of adiabaticity $\omega$. We plot coherence of oscillations for the same parameter range. These plots are made for parameter values $X^{\mathrm{eq}}=X^{\mathrm{ad}}=20$.
B Local winding number in combined bound and unbound chains
In this section we provide equations that are needed to define the local winding number in combined bound and unbound chains in the adiabatic phase. To define a local transition matrix, we choose a unit cell that consists of the $n$-th state in the bound chain and the $(n+1)$-th state in the unbound chain, together with transitions between these states as well as transitions to and from their right neighbors. Under periodic boundary conditions, the local transition matrix is
$\mathcal{W}^{n,\mathrm{tot}}(\chi)=\begin{pmatrix}(n+1)k(e^{-i\chi}-1)-f&(n+1)h\\ f&g_{1}(e^{i\chi}-1)+(n+2)k(e^{-i\chi}-1)-(n+1)h\end{pmatrix}.$ (S1)
We compute the combined local winding by plugging this transition matrix in Eq. 1 in the main text.
C Current balance follows from Kirchhoff’s laws
@@
In this section, we derive the current balance principle, which was used in the main text, from the Kirchhoff’s laws. To start, we define the probability current from the $i$-th to the $j$-th state as
$J_{ji}=\mathcal{W}_{ji}p_{i}-\mathcal{W}_{ij}p_{j}.$ (S2)
@@
$-J_{10,00}=J_{11,00}=J_{21,11}.$ (S5)
Summing up Eq. (S3) for $n$ from $1$ to $\infty$, and taking into account Eq. (S5), we derive that the vertical current summed over the network is zero
$\sum_{n=0}^{\infty}J_{(n+1)1,n0}=0.$ (S6)
D Analytical derivation of the domain wall in the adiabatic phase
In this section, we analytically derive the domain wall location for the combined local winding number in the adiabatic phase. For this, we use the fact that the winding number changes its sign when currents in the corresponding periodic network vanish [29]. To find this location, we track how vertical and horizontal currents in a local periodic network given by Eq. (S1) change as a function of protein number $n$. Assuming that the total probability in a local network splits into probability $p_{0}$ in the lower chain and $p_{1}$ in the upper chain, we can express the currents in the $(n-1)$-th local network as
$J_{\mathrm{vertical}}$ $=fp_{0}-nhp_{1},$ (S7)
$J_{\mathrm{horizontal}}$ $=J_{0}+J_{1}=-nkp_{0}+(g_{1}-(n+1)k)p_{1},$ (S8)
@@ D Analytical derivation of the domain wall in the adiabatic phase
$n^{*}_{\mathrm{tot}}=\left(-X^{\mathrm{eq}}+\sqrt{(X^{\mathrm{eq}})^{2}+4X^{\mathrm{eq}}(2X^{\mathrm{ad}}-1)}\right)/2.$ (S11)

### v3

@@ V Conclusion
Acknowledgments
@@
We thank Dexin Li and Erwin Frey for helpful insights into winding numbers and flux balance respectively. This work was supported by the NSF Center for Theoretical Biological Physics (PHY-2019745) and the NSF CAREER Award (DMR-2238667).
Appendix A Global spectrum
The gene switch is determined by a transition matrix $\mathcal{W}$ introduced in the main text. In this Appendix, we explain how we computed its spectral flow.
First, we take one link of the gene network $n^{*}\rightleftharpoons n^{*}+1$, and allow it to take on a complex value whose phase $\chi$ can be swept from $\pi$ to $\pi$, $\mathcal{W}_{(n^{*}+1)1,n^{*}1}(\chi)=\mathcal{W}_{(n^{*}+1)1,n^{*}1}(0)e^{i\chi}$ and $\mathcal{W}_{n^{*}1,(n^{*}+1)1}(\chi)=\mathcal{W}_{n^{*}1,(n^{*}+1)1}(0)e^{-i\chi}$. Note that this phase could be inserted on any rate. In our calculation, we choose
$n^{*}=\left(-X^{\mathrm{eq}}+\sqrt{(X^{\mathrm{eq}})^{2}+4X^{\mathrm{eq}}(2X^{\mathrm{ad}}-1)}\right)/2,$
which corresponds to the domain wall derived in Appendix E.
Then, we compute the spectrum of the periodic matrix $\mathcal{W}(\chi)$ using exact diagonalization for a range of $\chi\in[-\pi,\pi]$. In our calculation, the network consists of 60 states, 30 in each chain, bound and unbound. Each transition matrix is calculated for 21 value of $\chi$ equidistantly distributed between $\pi$ and $\pi$ values. Eigenvalues of the transition matrix $\mathcal{W}(\chi)$ flow continuously as the phase $\chi$ is swept between $\pi$ and $\pi$, thus forming a spectral flow.
Note that the spectra in Fig. 2(a) of the main text are zoomed in to distinguish fine details near the steady state, so not all calculated eigenvalues are visible in the figure. We plot the zoomed-out spectra in Fig. 5 for the same set of parameters.
Figure 5: Spectrum of the transition matrix $\mathcal{W}(\chi)$ with manually imposed periodic boundary conditions for a range of $\chi\in[-\pi,\pi]$ and three values of adiabaticity $\omega$. Other system parameters are set to $X^{\mathrm{ad}}=X^{\mathrm{eq}}=10$.
Appendix B Counting statistics interpretation of the field $\chi$
In this Appendix we provide a physical interpretation of the spectral flow introduced in Appendix A. It can be understood through the number of times that a typical trajectory crosses the link with the phase $\chi$, the so called counting statistics introduced in Ref. [36].To count the number of trajectory crossings, this formalism introduces a probability generating function
$Z(\chi)=\langle 1|e^{\mathcal{W}(\chi)T}|\bar{p}\rangle,$
where $T$ is time over which we observe the system, $\chi$ has the meaning of the counting field, and $\log Z(\chi)$ is the full counting statistics. This means that the derivative
$\frac{d}{d(i\chi)}\log Z(\chi)\big|_{\chi=0}$
gives the number of times that a typical trajectory crosses the chosen link over the time $T$. In our system, this corresponds to the number of clockwise loops that the system makes around the network. The corresponding current can be computed as the derivative of the first eigenvalue of the transition matrix [29]:
$J=\frac{d}{d(i\chi)}\lambda_{0}(\chi)\big|_{\chi=0}.$
@@ Appendix B Counting statistics interpretation of the field $\chi$
Figure 6: Exceptional point phase transition in the gene switch. We plot two closest to the steady state eigenvalues, and a scalar product of the corresponding eigenvectors as a function of adiabaticity $\omega$. We plot coherence of oscillations for the same parameter range. These plots are made for parameter values $X^{\mathrm{eq}}=X^{\mathrm{ad}}=20$.
Finally, we would like to compare this interpretation with a flux insertion method used to compute spectral flows in disordered quantum systems [37]. There, the phase $\chi$ is usually distributed evenly across all transitions of the entire network. In this case, phase $\chi$ can be interpreted as a magnetic flux inserted in the middle of a closed loop, affecting the phase of transition amplitudes. While the physical interpretation in quantum and stochastic systems is different, both methods enable computing spectral flows and studying topological properties of these systems.
Appendix C Exceptional point phase transition
In the main text, we discussed how three regimes of the gene switch are separated by exceptional point phase transitions [42, 41]. Here, we provide a more detailed analysis of these phase transitions.
Exceptional points are characterized by a simultaneous coalescence of eigenvalues and eigenvectors of a non-Hermitian matrix. We demonstrate that the two closest to the steady state eigenvalues of the transition matrix $\mathcal{W}$ go through an exceptional point. For this we plot in Fig. 6 how their real and imaginary parts change with adiabaticity $\omega=f/k$. We also plot a normalized scalar product of the corresponding eigenvectors, which turn to 1 when the two vectors become linearly dependent. This happens at the same values of $\omega$ at which the eigenvalues coalesce, proving that the system goes through exceptional points phase transitions.
Plotting in Fig. 6 the coherence of gene network oscillations for the same parameter range, we confirm that coherence changes from zero to non-zero values at exceptional point phase transitions.
Appendix D Current balance follows from Kirchhoff’s laws
@@
In this Appendix, we derive the current balance principle, which was used in the main text, from the Kirchhoff’s laws. To start, we define the probability current from the $i$-th to the $j$-th state as
$J_{ji}=\mathcal{W}_{ji}p_{i}-\mathcal{W}_{ij}p_{j}.$ (7)
@@
$J_{10,00}=J_{11,00}=J_{21,11}.$ (10)
Summing up Eq. (8) for $n$ from $1$ to $\infty$, and taking into account Eq. (10), we derive that the vertical current summed over the network is zero
$\sum_{n=0}^{\infty}J_{(n+1)1,n0}=0.$ (11)
Appendix E Analytical derivation of the domain wall in the adiabatic phase
In this Appendix, we analytically derive the domain wall location for the combined local winding number in the adiabatic phase. For this, we use the fact that the winding number changes its sign when currents in the corresponding periodic network vanish [29]. To find this location, we track how vertical and horizontal currents in a local periodic network given by Eq. (5) change as a function of protein number $n$. Assuming that the total probability in a local network splits into probability $p_{0}$ in the lower chain and $p_{1}$ in the upper chain, we can express the currents in the (n1)(n-1)-th local network as
$J_{\mathrm{vertical}}$ $=fp_{0}-nhp_{1},$ (12)
$J_{\mathrm{horizontal}}$ $=J_{0}+J_{1}=-nkp_{0}+(g_{1}-(n+1)k)p_{1},$ (13)
@@ Appendix E Analytical derivation of the domain wall in the adiabatic phase
$n^{*}_{\mathrm{tot}}=\left(-X^{\mathrm{eq}}+\sqrt{(X^{\mathrm{eq}})^{2}+4X^{\mathrm{eq}}(2X^{\mathrm{ad}}-1)}\right)/2.$ (16)